What is the minimum value of y = 2x^2-8x +11?
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

What is the minimum value of y = 2x^2-8x +11?
----
It's the vertex, on the Axis of Symmetry.
The AOS is x = -b/2a
x = 8/4 = 2
---
y = 2*4 - 8*2 + 11
= 3

Answer by ewatrrr(24785)

y = 2x^2-8x +11
y = 2(x^2-4x) + 11
y = 2(x-2)^2 - 2*4 + 11
y = 2(x-2)^2 + 3
minimum value of y = 3